The amount of an investment of P dollars for t years at a simple interest rate r is given by A= P+ Prt.
a) Rewrite this formula by factoring out the greatest common factor on the right hand side.
b) Find A if $8300 is invested for 3 years at a simple interest rate of 15%.

Answer by stanbon(75887)

The amount of an investment of P dollars for t years at a simple interest rate r is given by A= P+ Prt.
a) Rewrite this formula by factoring out the greatest common factor on the right hand side.
A = P(1+rt)
b) Find A if $8300 is invested for 3 years at a simple interest rate of 15%.
A = 8300(1+0.15*3)= $12,035
=================
Cheers,
Stan H.